OpenStudy (anonymous):

Which of the following is equivalent to 36/8? 4 1/4 4 1/2 4 2/3 4 3/4

OpenStudy (karatechopper):

well, reduce it to a mixed fraction!

OpenStudy (anonymous):

the second one 4 1/2

OpenStudy (anonymous):

36/8=9/2 and 4 1/2 = 4.5 which is the same as 9/2

OpenStudy (anonymous):

i did explain it!

OpenStudy (karatechopper):

o oops lol u posted it right when i posted mine:) srry

OpenStudy (anonymous):

yea i knw ppl won't get anything until we show them the steps :] its okay

OpenStudy (karatechopper):

haha yea